OpenStudy (anonymous):

Find two values of b that will make 64x^2+bx+49 a perfect square trinomial?

OpenStudy (lgbasallote):

hint: when you have \(ax^2 + bx +c\) then \[c = (\frac b2)^2\] \[\implies \sqrt c = \frac b2\] \[\implies 2\sqrt c = b\] so in this case, your c is 49 can you find b now?

OpenStudy (anonymous):

ok so what would my 2 values be?

OpenStudy (anonymous):

would it not be 7, 8 ?

OpenStudy (lgbasallote):

what is the square root of 49?

OpenStudy (anonymous):

7